# High Resolution UDF Meshing via Iterative Networks

### v1

@@ 4.2 Metrics
Figure 3: Meshing at different resolutions. While NSD-UDF [31] retrieves most of the surface well at a low resolution, it struggles at higher ones. In contrast, our method, recovers the surface well at all resolutions. We use Marching Cubes with both methods.
@@ 4.3 Comparing against other methods
In Tab. 2(a)2(b) and Fig. 4, we compare our method against state-of-the-art ones on four datasets, spanning from low to high resolutions. We divided the methods into two groups, those that rely on Marching Cubes and those that use Dual Contouring. Notably, DCUDF [12] and DualMesh-UDF [35] have been shown to be sensitive to the choice of their parameters. We report results obtained using both the default parameters and those that we manually tuned, denoted as “-T”. For technical reasons related to its implementation, the resolution of DualMesh-UDF and all methods that use it is +1+1 higher than the other methods. As it is a minor increase, we report them in the same table nonetheless. UNDC [5] failed to reconstruct meshes at high resolution due to the method’s VRAM requirements. DCUDF [12] failed to reconstruct some of the shapes in the experiments, producing unbound metrics. We discard such shapes from the reported results of DCUDF, which means that its numbers are not directly comparable to the other baselines.
Our method consistently outperforms all baselines at high resolutions ($256^{3}$ and $512^{3}$) on complex shapes such as cars, chairs, and planes, which are the most prone to suffer the UDF issues discussed at the beginning of Section 3 and depicted by Fig. 1. This is particularly visible in the three bottom rows of Fig. 4, where our method is the only one able to recover the front of the car, the legs of the chair and the cockpit of the plane. DCUDF [12] yields the smoothest meshes overall, at the cost of losing details and sometimes entire portions of the surface, as shown in Fig. 4. While previous comparisons [12, 31] have shown CAP-UDF [38] to be often be less accurate than other methods at low resolutions, we notice that the trend is reversed at higher resolutions. There, it retrieves surface portions missing from other baselines. MeshUDF [11] and NSD-UDF [31] can outperform it at high resolutions, but only when their filtering thresholds are removed, as shown in Tab. 3.
At lower resolutions the missing surface problem is less pronounced, and thus iterative refinement does not provide significant advantages. To illustrate this, we present examples of a 3D chair meshed at resolutions of $128^{3}$, $256^{3}$, and $512^{3}$ in Fig.˜3, comparing one of the most accurate methods, NSD-UDF [31], with our approach. As shown, at lower resolutions NSD-UDF reconstructs good surfaces, leaving less room for improvement, however it fails to recover many surface regions at $256^{3}$ and especially at $512^{3}$ resolution, whereas our method handles these cases more effectively. Similarly, our method does not show significant advantages on the MGN dataset, where the shapes are less complex and nearly all methods achieve high accuracy without missing large portions of the surface.
@@
GT CAP-UDF MeshUDF DCUDF-T NSD-UDF Ours UNDC DMUDF-T NSD-UDF Ours
Figure 4: Qualitative comparison with existing methods. Surface meshing results of neural UDFs with all methods at resolution of 512 (and 256 for MGN). UNDC failed at resolution 512 due to high GPU memory requirements. indicates that the method is combined with DMUDF.
@@ 4.4 Importance of Correct Thresholding
As described in Sec. 3.2, a correct filtering strategy is crucial to keeping the computational costs under control when meshing neural UDFs at high resolution. MeshUDF [11] and NSD-UDF [31] rely on filtering out high UDF values to speed up the computation, which can be far from lossless at high resolutions. Removing such thresholds can help the algorithms to better retrieve surfaces at high resolution, but this is insufficient by itself. Tab. 3 shows that there is a noticeable improvement at high resolutions, but the results are still not as good as those of our iterative pipeline. MeshUDF can become less accurate at resolution 256 due to artifacts introduced in regions of the space with high UDF values, which its heuristic is not designed to handle. In contrast, our method is robust to all such perturbations.
@@ 4.5 Mesh Evolution over Successive iterations
In Tab. 4 and Fig. 5, we show the evolution of the meshes over several iterations of our scheme. The meshes are progressively refined and the missing regions filled, with the Chamfer Distance decreasing in most cases, especially at higher resolutions and in complicated shapes such as cars, chairs and planes. The only partial exception to this trend occurs in the MGN dataset, where the mesh is already very accurate after the first iteration and remains so during the following ones at high resolutions. At a lower resolution, such as 64, we see the opposite behavior, especially for simple shapes like garments. The problems described in Fig. 1 do not arise, lessening the usefulness of an iterative scheme. We also observe that our pipeline often produces more accurate meshes than the baselines already during the first iteration, as shown in Fig. 1 and by comparing results in Tab. 4 and 3. This is mainly due to the more robust training of our network compared to NSD-UDF [31] and thanks to the asbsence of heavy thresholds during meshing.
@@ 5 Limitations
For simpler shapes, such as garments, methods specifically designed to prioritize topology over accuracy, such as MeshUDF, may sometimes be more suitable. Moreover, while allowing high resolution surface extraction, our method can still produce artifacts in the form of irregular boundaries and small holes, as visible in Figs.˜4 and 5. Post-processing algorithms can help with this, but future work should focus on improving in these aspects.
@@ 6 Conclusion
A.1 Mean results and standard deviation
Table A.1: Triangulating Neural Unsigned Fields using Marching Cubes-based method. L2 Chamfer Distance $\times 10^{-5}$ with 2M sample points (CD), F1 score (F1) and Image Consistency (IC) are reported at varying grid resolutions. The best results are in bold. DCUDF failed to mesh some of the shapes, but its numbers are reported nonetheless. *Resolution is halved for experiments with MGN due to the lower complexity of the shapes.
(a) Mean results.
MGN* [1] ShapeNet cars [3] ShapeNet chairs [3] ShapeNet planes [3]
128-MC CAP-UDF [38] 19.7 68.2 78.6 71.7 49.5 81.8 518 51.3 68.9 19.5 68.1 79.0
MeshUDF [11] 2.79 82.0 93.3 11.9 57.9 88.7 21.8 68.0 88.7 19.4 73.5 81.0
DCUDF [12] 14100 3.40 3.84 8000 10.6 13.7 22700 14.4 16.8 2900 28.9 24.3
DCUDF-T [12] 116 3.04 86.6 66.5 55.4 86.8 2720 65.2 78.2 902 69.4 78.8
NSD-UDF + MC [31] 1.55 82.9 94.1 9.40 59.9 88.5 17.7 68.7 88.5 4.72 78.0 84.5
Ours + MC 2.26 80.8 93.4 6.79 59.5 88.9 7.27 67.8 89.5 3.49 77.6 84.8
256-MC CAP-UDF [38] 3.23 85.4 91.2 48.3 58.8 86.2 223 65.5 78.5 8.95 83.2 85.0
MeshUDF [11] 1.16 88.2 94.5 17.1 61.1 88.2 67.1 69.8 85.4 4.83 85.2 85.1
DCUDF [12] 18200 5.17 4.12 1090 50.6 75.1 7900 50.7 55.3 278 82.4 80.2
DCUDF-T [12] 10.3 86.1 94.8 1080 50.3 75.0 7840 50.7 55.3 797 80.1 77.9
NSD-UDF + MC [31] 0.973 88.7 94.8 14.7 61.2 87.6 53.7 69.8 85.4 3.81 86.9 85.2
Ours + MC 1.02 87.6 94.4 7.26 63.1 89.0 10.6 70.6 89.3 2.64 88.2 86.6
512-MC CAP-UDF [38] 2.14 89.1 94.1 48.6 60.2 86.4 202 67.9 79.3 8.94 87.3 85.5
MeshUDF [11] 1.18 89.3 94.4 136 54.4 78.8 799 57.5 64.2 21.2 87.4 83.4
DCUDF [12] 49.4 85.8 88.8 478 52.9 80.2 7510 55.7 63.8 199 84.1 81.0
DCUDF-T [12] 32.8 85.9 88.9 478 52.9 80.2 7380 55.9 64.2 284 83.2 79.9
NSD-UDF + MC [31] 1.08 89.5 94.4 80.8 56.8 82.0 394 62.9 71.2 13.4 88.2 83.8
Ours + MC 0.880 89.3 94.4 12.3 63.4 88.2 36.6 71.6 86.5 3.31 90.1 86.5
(b) Standard deviation for each metric, computed across the dataset.
MGN* [1] ShapeNet cars [3] ShapeNet chairs [3] ShapeNet planes [3]
Res. Method CD F1 IC CD F1 IC CD F1 IC CD F1 IC
128-MC CAP-UDF [38] 12.1 7.83 3.46 51.9 12.2 5.69 649 15.4 14.7 21.3 5.63 3.45
MeshUDF [11] 1.28 7.9 2.34 5.21 12.1 2.22 30.3 13.9 5.58 27.4 5.54 2.56
DCUDF [12] 561 0.901 1.08 5560 4.81 5.74 18800 8.78 16.0 3010 7.84 7.45
DCUDF-T [12] 169 1.30 3.97 57.6 11.4 4.28 5020 14.0 17.6 1400 6.72 6.79
NSD-UDF + MC [31] 0.817 7.87 1.91 5.24 12.2 2.33 28.0 13.7 5.64 2.87 6.10 2.28
Ours + MC 0.872 7.75 2.09 3.16 11.7 1.92 8.55 13.4 4.93 1.49 6.02 1.89
256-MC CAP-UDF [38] 4.83 7.23 2.37 40.8 12.6 4.46 220 15.9 12.7 7.44 5.32 3.03
MeshUDF [11] 0.733 6.78 1.56 11.7 11.6 2.64 74.9 14.3 8.51 4.18 5.51 2.41
DCUDF [12] 12700 2.22 1.86 3250 14.0 15.2 9500 17.7 19.9 664 8.45 6.52
DCUDF-T [12] 12.5 7.00 1.68 3250 14.0 15.1 9440 17.6 19.9 1440 8.70 7.70
NSD-UDF + MC [31] 0.636 6.78 1.45 10.3 11.8 2.79 68.9 14.4 8.36 3.04 5.80 3.13
Ours + MC 0.557 6.55 1.46 3.91 10.8 1.95 13.0 14.0 5.12 2.37 4.37 2.06
512-MC CAP-UDF [38] 3.97 6.78 1.80 40.6 12.4 4.19 217 13.8 12.2 7.65 4.03 3.20
MeshUDF [11] 1.30 6.63 1.42 115 12.9 8.34 964 16.3 17.8 20.0 5.30 4.67
DCUDF [12] 362 10.2 7.60 989 14.7 12.8 9890 19.0 22.0 537 8.35 4.99
DCUDF-T [12] 219 9.98 7.33 989 14.7 12.8 9950 19.0 22.5 774 8.53 5.93
NSD-UDF + MC [31] 1.09 6.66 1.48 65.7 12.8 6.40 442 15.1 15.8 11.3 5.08 4.38
Ours + MC 0.563 6.49 1.27 8.41 11.2 2.48 48.5 12.7 7.55 2.83 4.08 2.73
Table A.2: Triangulating Neural Unsigned Fields using Dual Contouring-based methods. L2 Chamfer Distance $\times 10^{-5}$ with 2M sample points (CD), F1 score (F1) and Image Consistency (IC) are reported at varying grid resolutions. The best results are in bold. UNDC failed at resolution 512 due to high GPU memory requirements. *Resolution is halved for experiments with MGN due to the lower complexity of the shapes.
(a) Mean results.
MGN* [1] ShapeNet cars [3] ShapeNet chairs [3] ShapeNet planes [3]
128-DC UNDC [5] 1.26 86.1 93.8 17.1 60.8 86.0 91.3 66.8 78.8 3.82 81.8 85.7
DualMesh-UDF [35] 1120 61.1 60.9 1600 33.0 42.8 12400 15.0 13.0 341 72.4 72.3
DualMesh-UDF-T [35] 0.939 88.7 94.9 7.45 62.6 89.4 19.1 71.4 88.7 2.91 83.3 87.3
NSD-UDF + DualMesh-UDF [31] 0.901 89.2 94.5 8.60 63.9 89.1 17.3 70.2 88.3 3.12 85.9 86.5
256-DC UNDC [5] 1.32 87.9 91.2 122 49.2 69.0 598 50.9 55.9 21.4 81.9 78.7
DualMesh-UDF [35] 993 60.0 59.0 1440 33.3 42.6 11900 15.2 13.0 158 76.1 72.5
DualMesh-UDF-T [35] 0.899 89.7 94.7 14.3 61.8 86.6 62.2 69.0 82.4 3.26 87.8 86.8
NSD-UDF + DualMesh-UDF [31] 0.838 89.5 94.3 15.6 61.8 87.0 56.0 68.3 82.4 3.72 88.3 85.4
512-DC UNDC [5] 6.86 83.0 81.2 - - - - - - - - -
DualMesh-UDF [35] 948 58.4 57.1 1560 32.2 41.1 12400 14.7 12.2 172 75.6 70.6
DualMesh-UDF-T [35] 1.25 88.8 92.7 51.7 55.4 77.9 211 60.6 67.2 7.68 87.7 83.7
NSD-UDF + DualMesh-UDF [31] 1.13 88.2 91.9 86.3 54.3 77.8 412 58.4 63.9 13.8 86.7 82.1
(b) Standard deviation for each metric, computed across the dataset.
MGN* [1] ShapeNet cars [3] ShapeNet chairs [3] ShapeNet planes [3]
128-DC UNDC [5] 0.651 6.80 1.43 11.6 12.0 3.26 94.4 15.5 11.2 3.33 5.77 2.69
DualMesh-UDF [35] 2470 25.6 25.4 2090 12.0 13.2 12800 12.0 11.3 995 10.8 12.2
DualMesh-UDF-T [35] 0.561 6.51 1.36 4.17 10.5 2.10 29.3 13.5 5.94 2.58 5.34 1.94
NSD-UDF + DualMesh-UDF [31] 0.546 6.48 1.50 5.14 11.4 2.21 27.2 13.3 5.55 2.76 4.96 2.33
256-DC UNDC [5] 1.25 6.80 2.41 97.7 11.9 9.59 716 17.6 16.1 21.1 6.89 5.63
DualMesh-UDF [35] 2280 25.9 25.5 1390 11.6 12.8 12900 11.9 10.9 241 9.85 11.8
DualMesh-UDF-T [35] 0.673 6.49 1.31 9.94 10.7 3.12 72.3 13.6 8.97 2.96 3.96 2.39
NSD-UDF + DualMesh-UDF [31] 0.565 6.80 1.37 11.1 11.8 3.10 71.2 13.1 8.33 3.07 4.44 3.16
512-DC UNDC [5] 11.3 9.36 6.95 - - - - - - - - -
DualMesh-UDF [35] 2080 26.0 25.6 1960 11.4 12.5 14500 11.6 10.3 267 9.88 11.7
DualMesh-UDF-T [35] 1.36 7.14 2.51 38.5 11.5 5.96 221 13.6 14.3 6.87 4.94 3.93
NSD-UDF + DualMesh-UDF [31] 1.18 7.48 2.92 69.7 12.6 7.41 467 14.9 15.5 11.4 5.2 4.83
For the sake of completeness, we report in Tabs.˜1(a), 1(b), 2(a) and 2(b) the mean results on the datasets and methods used in the main paper, alongside their respective standard deviations. We observe that the trends are consistent with the median results shown in the main paper, with our method outperforming all baselines at high resolutions on complex shapes, and competing closely with existing methods at lower resolutions and on simpler shapes. The standard deviations computed on our method are also consistently lower than the baselines at high resolution, while remaining competitive at lower resolutions, showing that our method is also more robust to shape variations compared to existing baselines. As in the main tables, notice that UNDC [5] failed to reconstruct meshes at high resolution due to the method’s VRAM requirements. DCUDF [12] failed to reconstruct some of the shapes in the experiments, producing unbound metrics. We discard such shapes from the reported results of DCUDF, which means that its numbers are not directly comparable to the other baselines.
A.2 Implementation details

### v2

@@ 4.2 Metrics
GT CAP-UDF MeshUDF DCUDF-T DCUDF-T-nocut NSD-UDF Ours UNDC DMUDF-T NSD-UDF Ours
Figure 3: Qualitative comparison with existing methods (auto-decoders). Surface meshing results of neural UDFs with all methods at resolution of 512 (and 256 for MGN). UNDC failed at resolution 512 due to high GPU memory requirements. indicates that the method is combined with DMUDF.
Table 3: Neural Unsigned Distance Fields from point clouds. L2 Mesh Chamfer Distance $\times 10^{-5}$ with 2M sample points (CD), F1 score (F1) and Image Consistency (IC) are reported at varying grid resolutions. Median scores are reported for cars; mean for scenes due to the low number of samples. The best results are in bold.
CAP-L scenes [Zhou24] CAP-L cars [Zhou24] DiffUDF cars [Fainstein24]
Res. Method CD $\downarrow$ F1 $\uparrow$ IC $\uparrow$ CD $\downarrow$ F1 $\uparrow$ IC $\uparrow$ CD $\downarrow$ F1 $\uparrow$ IC $\uparrow$
128 CAP-UDF [Zhou24] 4.27 68.0 83.8 10.7 44.5 85.8 6.71 58.0 86.3
MeshUDF [Guillard22b] 4.40 68.2 84.5 9.83 45.4 85.7 9.02 59.7 86.7
DCUDF-T [Hou23a] 279 59.7 84.0 45.5 41.4 85.5 32.7 49.6 89.6
DCUDF-T-nocut [Hou23a] - - - 8.47 47.2 88.3 12.5 56.9 90.2
NSD-UDF + MC [Stella24] 3.34 69.8 86.7 8.14 47.1 86.4 4.18 65.7 87.7
Ours + MC 3.44 69.3 85.3 7.03 48.3 86.8 5.46 62.3 87.4
256 CAP-UDF [Zhou24] 3.65 70.1 86.1 7.51 48.0 86.6 3.72 67.9 86.7
MeshUDF [Guillard22b] 3.42 69.6 86.3 8.65 47.5 86.1 5.18 65.9 87.2
DCUDF-T [Hou23a] 4.75 69.9 84.6 31.1 45.5 85.7 25.5 58.3 86.4
DCUDF-T-nocut [Hou23a] - - - 9.58 47.9 86.0 5.56 69.5 87.7
NSD-UDF + MC [Stella24] 3.30 70.3 86.3 8.76 47.9 86.3 3.63 70.0 86.9
Ours + MC 3.07 71.4 86.6 7.22 50.1 87.1 3.45 68.9 88.3
512 CAP-UDF [Zhou24] 3.57 70.5 86.1 8.11 48.1 86.4 4.60 67.1 84.4
MeshUDF [Guillard22b] 4.49 69.9 84.5 8.46 47.6 85.9 3.83 67.4 86.4
DCUDF-T [Hou23a] 140 68.9 83.3 31.5 45.4 85.8 17.2 60.1 84.3
DCUDF-T-nocut [Hou23a] - - - 11.8 46.3 84.8 5.53 69.0 87.1
NSD-UDF + MC [Stella24] 3.72 70.0 84.0 9.85 47.3 85.9 5.63 66.0 82.0
Ours + MC 3.08 71.8 86.8 7.82 49.8 86.6 3.03 71.7 88.1
GT CAP-UDF MeshUDF DCUDF-T DCUDF-T-nocut NSD-UDF Ours DMUDF NSD-UDF Ours
Figure 4: Qualitative comparison with existing methods (reconstruction from point clouds). Surface meshing results of neural UDFs at resolution of 512. Top: CAP-L 3D scenes. Middle: CAP-L car. Bottom: DiffUDF cars. indicates that the method is combined with DMUDF.
@@
Figure 5: Meshing at different resolutions. While NSD-UDF [Stella24] retrieves most of the surface well at a low resolution, it struggles at higher ones. In contrast, our method, recovers the surface well at all resolutions. We use Marching Cubes with both methods.
@@ 4.3 Comparing against other methods
In Tab. 2(a)2(b) and Fig. 3, we compare our method against state-of-the-art ones using four auto-decoders on four datasets, spanning from low to high resolutions. Our method consistently outperforms all baselines at high resolutions ($256^{3}$ and $512^{3}$) on complex shapes such as cars, chairs, and planes, which are the most prone to suffer from the UDF issues discussed at the beginning of Section 3 and depicted by Fig. 1. This is particularly visible in the three bottom rows of Fig. 3, where our method is the only one able to recover the front of the car, the legs of the chair and the cockpit of the plane. DCUDF [Hou23a] yields the smoothest meshes overall, at the cost of losing details and sometimes entire portions of the surface, as shown in Fig. 3, unless the cut operation is avoided, resulting in double surfaces. While previous comparisons [Hou23a, Stella24] have shown CAP-UDF [Zhou24] to be often be less accurate than other methods at low resolutions, we notice that the trend is reversed at higher resolutions. There, it retrieves surface portions missing from other baselines. MeshUDF [Guillard22b] and NSD-UDF [Stella24] can outperform it at high resolutions, but only when their filtering thresholds are removed, as shown in Tab. 4.
At lower resolutions the missing surface problem is less pronounced, and thus iterative refinement does not provide significant advantages. To illustrate this, we present examples of a 3D chair meshed at resolutions of $128^{3}$, $256^{3}$, and $512^{3}$ in Fig.˜5, comparing one of the most accurate methods, NSD-UDF [Stella24], with our approach. As shown, at lower resolutions NSD-UDF reconstructs good surfaces, leaving less room for improvement, however it fails to recover many surface regions at $256^{3}$ and especially at $512^{3}$ resolution, whereas our method handles these cases more effectively. Similarly, our method does not show significant advantages on the MGN dataset, where the shapes are less complex and nearly all methods achieve high accuracy without missing large portions of the surface.
@@
In Tab. 3 and Fig. 4, we show additional experiments using two single-shape neural UDF architectures trained on point clouds: CAP-L [Zhou24] and DiffUDF [Fainstein24]. For 3D scenes we used the $100~points/m^{2}$ setting from the CAP-UDF [Zhou24] paper, using the provided data. For cars, we sample 10k noise-free points per shape following the CAP-UDF protocol. Both architectures are trained using default parameters from their respective repositories. The results follow similar trends as in the auto-decoder experiments: our method achieves higher accuracy than the baselines, especially at high resolutions. Additional DC-based results and experimental details are included in the supplementary material.
@@ 4.4 Importance of Correct Thresholding
As described in Sec. 3.2, a correct filtering strategy is crucial to keeping the computational costs under control when meshing neural UDFs at high resolution. MeshUDF [Guillard22b] and NSD-UDF [Stella24] rely on filtering out high UDF values to speed up the computation, which can be far from lossless at high resolutions. Removing such thresholds can help the algorithms to better retrieve surfaces at high resolution, but this is insufficient by itself. Tab. 4, computed on the auto-decoder experiments, shows that there is a noticeable improvement at high resolutions, but the results are still not as good as those of our iterative pipeline. MeshUDF can become less accurate at resolution 256 due to artifacts introduced in regions of the space with high UDF values, which its heuristic is not designed to handle. In contrast, our method is robust to all such perturbations.
@@ 4.5 Mesh Evolution over Successive iterations
In Tab. 5 and Fig. 6, we show the evolution of the meshes over several iterations of our scheme on auto-decoder-based UDFs. The meshes are progressively refined and the missing regions filled, with the Chamfer Distance decreasing in most cases, especially at higher resolutions and in complicated shapes such as cars, chairs and planes. The only partial exception to this trend occurs in the MGN dataset, where the mesh is already very accurate after the first iteration and remains so during the following ones at high resolutions. At a lower resolution, such as 64, we see the opposite behavior, especially for simple shapes like garments. The problems described in Fig. 1 do not arise, lessening the usefulness of an iterative scheme. We also observe that our pipeline often produces more accurate meshes than the baselines already during the first iteration, as shown in Fig. 1 and by comparing results in Tab. 5 and 4. This is mainly due to the more robust training of our network compared to NSD-UDF [Stella24] and thanks to the asbsence of heavy thresholds during meshing.
@@ 5 Limitations
For simpler shapes, such as garments, methods specifically designed to prioritize topology over accuracy, such as MeshUDF, may sometimes be more suitable. Moreover, while allowing high resolution surface extraction, our method can still produce artifacts in the form of irregular boundaries and small holes, as visible in Figs.˜3 and 6. Post-processing algorithms can help with this, but future work should focus on improving in these aspects.
@@ 6 Conclusion
7 Acknowledgements
This work was funded in part by the Swiss National Science Foundation.
A.1 Complete quantitative results
For the sake of completeness, we report in Tabs.˜3(a), 3(b), 4(a), 4(b), 5(a), 5(b), 6(a) and 6(b) the mean results on the datasets and methods used in the main paper, alongside their respective standard deviations. We also report the results of DC-based methods on point cloud reconstruction from CAP-L [Zhou24] and DiffUDF [Fainstein24] in Tab.˜A.2. We observe that the trends are consistent with the median results shown in the main paper, with our method outperforming all baselines at high resolutions on complex shapes, and competing closely with existing methods at lower resolutions and on simpler shapes. The standard deviations computed on our method are also consistently lower than the baselines at high resolution, while remaining competitive at lower resolutions, showing that our method is also more robust to shape variations compared to existing baselines. As in the main tables, notice that UNDC [Chen22b] failed to reconstruct meshes at high resolution due to the method’s VRAM requirements. DCUDF [Hou23a] failed to reconstruct some of the shapes in the experiments, producing unbound metrics. We discard such shapes from the reported results of DCUDF, which means that its numbers are not directly comparable to the other baselines.
A.2 Implementation details and other ablation studies